OpenStudy (anonymous):

y=x-3 whats the slope ?

OpenStudy (nurali):

The linear equation y = mx + b represents a straight line with slope = m and y-intercept = b y = x - 3 m = 1 so the slope is 1 b = -3 so the y-intercept is -3